OpenStudy (anonymous):

im practicing for a placement test and im having trouble with fractions add and subtraction like Example:15 5/16-6 3/8 how do you get the answer

OpenStudy (anonymous):

i think the best way to add these mixed fractions is to turn them into improper fractions, then find the common dneminator and add them like any other fractions

OpenStudy (anonymous):

i think lcm makes u eaisy.

OpenStudy (radar):

I like to solve piece-work style. 15-6=9 5/16-3/8=5/16-6/16=-1/16 9-1/16=8 15/16

OpenStudy (chaise):

Convert from improper fractions you multiply the denominator by the number out front then add the numerator. The entire thing is divided by the denominator. You then need to get equal denominators to add or subtract the two fractions. 15 and 5/16 = 16x15+5/16 = 245/16 16 and 6/16 = 16x16+5/16 261/16 261/16 - 245/16 = -1

OpenStudy (chaise):

Oh, my bad, It's the other way round. 245/16 - 261/16 = -1 Sorry.

OpenStudy (anonymous):

thank u